the flower garden has the shape of a right triangle. 17ft of a perennial border forms the hypotenuse of the triangle, and one leg is 7 ft longer than the other leg. find the length of the legs
Answer by Cromlix(4381)

Using the Pythagorean formula
Base leg = x
Vertical leg = x + 7
Hypotenuse = 17
x^2 + (x + 7)^2 = 17^2
x^2 + x^2 + 14x + 49 = 289
Collect like terms;
2x^2 + 14x - 240 = 0
Divide by 2
x^2 + 7x - 120 = 0
(x - 8)(x + 15) = 0
x + 15 = 0
x = -15 (no answer x < 0)
x - 8 = 0
x = 8
Size of legs:
Base leg = 8 ft
Vertical leg = 15 ft
Hope this helps.
:-)